Grade 12Algebra

Ritu ate (3/5) part of an apple and the remaining apple was eaten by her brother Somu. How much part of the apple did Somu eat? Who had the larger share? By how much?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question, it is given that,
Part of apple eaten by Ritu is = (3/5)
Part of apple eaten by Somu is = 1–Part of apple eaten by Ritu
= 1–(3/5)
The LCM of 1, 5 = 5
Now, let us change each of the given fraction into an equivalent fraction having 10 as the denominator.
= [(1/1) × (5/5)]–[(3/5) × (1/1)]
= (5/5)–(3/5)
= (5 - 3)/5
= 2/5
∴Part of apple eaten by Somu is (2/5)
So, (3/5) > (2/5)
hence, Ritu ate larger size of apple.
Now, the difference between the 32 shares = (3/5)–(2/5)
= (3 - 2)/5
= 1/5
Thus, Ritu’s share is larger than share of Somu by (1/5)

Thanks